If alpha and beta are the zeros of p(x)= 3x^2+5x-2 then form a quadratic polynomial whose zeros are 2alpha and 2 beta

Grade:10

1 Answers

sonika p
6 years ago
given alpha , beta are zeroes of p(x) = 3x2 + 5x -2thus, alpha + beta = -b/a = -5/3alpha*beta = c/a = -2/3if zeroes are 2alpha, 2beta2alpha+2beta=2(alpha+beta)= -10/32alpha*2beta = 4*alpha*beta = -8/3the new polynomial is given by q(x) = x^2 +(10/3)x -(8/3) or 3x^2+10x-8